# High-entropy Advantage in Neural Networks’ Generalizability

### v1

@@ Discussion
In this paper, we investigated the relationship between the entropy and generalizability of neural networks (NNs). We preformed experiments in four distinctive machine learning tasks, covering arithmetic questions, real world tabular data, computer vision, and language modeling. WLMC and WLMD methods are used to sample the NNs’ entropy landscapes for both classification ($S(L_{train},A_{test})$) and regression tasks ($S(L_{train},L_{test})$). For classification tasks, a smoothed test accuracy is employed for WLMD-simulated networks to guarantee the differentiability. At high train loss ($L_{train}$), the max-entropy states achieve random-guess accuracies, supporting the validity of the Wand-Landau-based algorithm in sampling the entropy of NNs. For networks with low train loss value, the max-entropy states generally achieve higher accuracy than the SGD-trained states. For regression tasks, the max-entropy states have lower or at least comparable test loss at all levels of train loss.
@@
Our contributions are two-fold. First, such a high-entropy advantage may be widely present in different neural networks architectures and learning tasks. If so, our work shed lights on a new approach to enhancing model performance without changing its architecture: developing optimizers that can effectively find NNs’ equilibrium (max entropy) states. This could be achieved by methods like low-temperature Newtonian molecular dynamics simulations or simulated annealing. Second, the high-entropy advantage also hints on a potential theoretical explanation for the generalizability of overparametrized models after SGD training. The SGD training algorithm minimizes the loss function using an estimated the gradient of the loss function. The inaccuracy involved in such estimation may bias the trained NN toward higher-entropy minima, which possess higher generalizability.
@@
While there are many studies focusing on using machine learning models to solve problems in physics [41, 42, 43, 44, 45], the reverse directions retains much less attention. However, the foundation concepts of machine learning, especially for neural networks, are deeply rooted in physics. This historic connection is exemplified by the pioneering work including the Hopfield networks [46, 47] and the Boltzmann machine [48, 49], which is draw directly from principles of statistical mechanics. We believe our work (re)highlights the potential of treating NNs as physical systems, promoting possibilities to apply well-established statistical physics methods to aid machine learning research. It would be intriguing to explore how techniques like simulated annealing or parallel tempering could accelerate the searching process of NNs’ equilibrium state. While a few studies have explored this direction before[50, 51], the existence of the high-entropy advantage in NNs underscores the great potential of further research.
@@ Methods
Wand-Landau Monte Carlo for Neural Networks
@@
WLMC is employed for the Spiral Classification Dataset and our results are averaged over 6 independently generated training and testing datasets. For each dataset, 10,000 stages simulations has been performed, where each stage contains $3.2\times 10^{6}$ steps. In each step, we randomly change one parameter in the network and then evaluate both train loss and test accuracy. Therefore, WLMC algorithm has a time complexity of $O(N)$ to update one trainable parameter, where $N$ is the total number of parameters in the network. We also employed an adapted step size during the simulation by calculating the accepting rate for every 1000 steps. The step size is increased by $10\%$ if the rate is higher than 0.7 and is decreased by $10\%$ when it is lower than 0.3. The ’modification factor’ in original Wang-Landau method[52] is set to be $f=\exp[5/(n+10)]$ for the nth stage.
@@
Parameters in the neural networks ranges from $\infty$- ∞ to $\infty$, when no constraint is added. Therefore, the volume of a parameter space is infinite, even for a tiny neural network. On the other hand, to make the thermodynamic quantities valid we must prescribe finite ranges for the parameters. According to early-established experience in neural network training, to alleviate the exploding-gradient and diminishing-gradient problems, the weight parameters should be on the order of $1/\sqrt{W}$ [53]. Thus, we limit weight parameters to the range of $[-2/\sqrt{W},2/\sqrt{W}]$ for our WLMC simulation.
Wand-Landau Molecular Dynamics for Neural Networks
@@
When sampling the entropy landscape of practical neural networks with tens of thousands parameters, WLMC becomes too expensive as it needs to perform the entire forward propagation to update one parameter. Therefore, we employed the WLMD method proposed before[54, 55], which has a time complexity of $O(N)$ to update $N$ parameters, where N is the total number of trainable parameters in the network. In addition, a Langevin thermostat is implemented to achieve the smooth temperature control during the simulation. The weight parameters are constrained by the range of $[-2/\sqrt{W},2/\sqrt{W}]$ and a reflecting boundary is applied. We also find that the range of the weights does not qualitatively change our results within a reasonable range.
@@
Unlike WLMC, the WLMD algorithm can only directly calculate the entropy as a function of differentiable variables. In order to calculate $S$ as a function of $A_{test}$, we designed a “smoothed” approximation of the test accuracy:
$A^{s}_{test}=S(\alpha\times\delta l),$ (2)
@@
where $S(x)=1/[1+\exp(-x)]$ is the sigmoid function, $\alpha$ is a hyperparameter, and $\delta l$ is the difference between the logit of the correct answer and the highest logit of incorrect answers. In the $\alpha\to\infty$ limit, we would have $A^{s}_{test}=A_{test}$. When $\alpha$ is finite, $A^{s}_{test}$ becomes a differentiable approximation of $A_{test}$. In practice we choose $\alpha=5$. This leads to a smoothing error, $|A^{s}_{test}-A_{test}|$, that is always smaller than $1.5\%$ except for the initial $10^{7}$ MD time steps. The initial part of the simulation have significantly higher smoothing errors (as much as $30\%$), but that does not affect the correctness of our final results since WLMD has the capability of starting from an incorrect entropy landscape and gradually converging to the correct entropy landscape.
@@
where $c_{i}=3000$ is a constant and $H(x)=\begin{cases}0,&x<0\\ 1,&x\geq 0\end{cases}$ is the Heaviside step function. Parameters $x_{{max}}$ and $y_{{max}}$ are system-dependent and listed in the supplementary material. By starting with such an $S_{{init}}$, we make the algorithm concentrate its exploration effort in the $x<x_{{max}}$ and $y<y_{{max}}$ regime, improving efficiency.
@@ Acknowledgements
To verify the correctness of the entropy diagram calculated from Wang-Landau MD simulations, we performed regular (Newtonian) MD of the MNIST system. We observed that as the system evolves and approaches equilibrium, the test accuracy indeed increases and approaches the high-entropy value.
The MD simulation was performed at temperature $k_{B}T=0.005$ with Langevin thermostat, and the system eventually equilibrates at potential energy $E_{p}=L_{\mbox{train}}\approx 0.04$. From the entropy diagram (Fig. 2b), at this $L_{\mbox{train}}$ the equilibrium test accuracy is around 0.76, while SGD training achieved a test accuracy of 0.68. Our MD results are presented in Fig. 4. The test accuracy is indeed slowly increasing and approaching 0.76, but the train loss is also slowly decreasing. Since the test accuracy can increase due to either decreasing train loss or increasing entropy (as time increases), we further performed a separation-of-variables study by plotting the test accuracy for only the time steps with the train loss within a small window of $0.0410.04<L_{\mbox{train}}<$ 0.041. For this case, the test accuracy still increases over time, which is a strong evidence that the test accuracy is indeed positively correlated with entropy.
Figure 4: (left) The train loss (energy) and test accuracy versus time steps (epoches) during Newtonian molecular dynamics simulations of the MNIST system. We averaged over 4 independent instances of simulations to reduce noise. (right) The test accuracy at time steps for which the train loss is in a small range of $0.0410.04<L_{\mbox{train}}<$ 0.041. Note that the length of error bars indicate the standard deviations of individual instances of neural networks, not the error estimation of each data point.
@@
Spiral Dataset
We generated the spiral dataset using the following equations:
@@
where $x_{1i}$ and $x_{2i}$ are the horizontal and vertical coordinates of a point, $r_{i}$ is uniformly distributed between 1 and 5, and $y_{i}$ is the label (i.e., color) of a point. Both $N_{1i}$ and $N_{2i}$ are Gaussian random noises with mean 0 and standard deviation 0.1.
Figure 5: (left) The spiral dataset. An independent test set was then generated following exactly the same procedure but with a different random seed. (right) Same left, except for the spiral regression dataset with 250 points.
@@
We generated the spiral regression datasets by first generating random coordinates, $-5<x_{1i}<5$ and $-5<x_{2i}<5$, for $N=50$, 100, or 250 points. We then calculate the label (color) of each point using the formula
@@
where $r_{i}$ and $\theta_{i}$ are the polar coordinates of the $i$th point.
@@
The MNIST handwritten digits dataset is downloaded using the torchvision Python package. We randomly selected 250 images for the training dataset and 250 images for the test dataset. We greatly reduced size of the dataset in order to increase the difficulty of machine learning and decrease the simulation time.
@@
The Polymer SMILES dataset is collected from the TransPolymer work[33], which is the foundation chemical language model we used in this study. The dataset is originally built by Ramprasad and coauthors, which consists of $561$ polymer SMILES strings and their bulk electric band gap energy (eV) [56, 36]. TransPolymer takes SMILES strings directly into the encoder, which is pretrained over $5$ millions polymer SMILES string via mask language modeling[33]. This pretrained encoder can generate an embedding vector with a dimension of $768$ for each polymer SMILES, which is then fed into the regressor head for downstream task learning. For sampling efficiency, we fixed the encoder of TransPolymer and only finetuning the regressor head, which is also commonly employed strategy for language model fine-tuning [37, 38, 33]. We follow the same protocol reported in the original TransPolymer work and randomly selected $80\%$ data for train set and the remaining $20\%$ data is used as test set.
@@
To benchmark the max-entropy state, we trained the corresponding neural networks for each task via the SGD optimizer. For a given network architecture, two hyperparameters have been tuned including learning rate and batch size. Best hyperparamters for different neural networks can be found in Table 1 to 5. Note for Kaggle House Price Prediction task, three different batch sizes have been employed for all neural networks to collect enough data for the full range of train loss. To make sure we are comparing max-entropy state with the best-available SGD training state, we always select the lowest test loss value for SGD training if there are more than one network falls into the bin of a certain train loss level. After determining the best training hyperparameters, we trained 100 neural networks and collected the train loss-test loss trajectories (regression task) or the train loss- test accuracy (classification task) during the entire training process of each task. These trajectories are then used to calculate the SGD training curves presented in Figure 1 to 3.
@@
To confirm the existence of entropy advantage in deeper neural networks with more complex architectures, we have constructed a 10-layer ResNet [29]. The neural network consists of an initial convolutional layer with 6 channels, and then 4 residual blocks with 6, 12, 12, and 12 channels each, and a final fully-connected layer. The original ResNet architecture contains batch-normalization operations, which breaks the rigorous analogy between training losses and physical energy landscapes. Therefore, we replaced all of them with layer normalization operations. Our training and test datasets each contain 2500 randomly chosen images of either a deer or a frog from the CIFAR-10 dataset. Due to the significantly higher computational cost, we only ran WLMD for $3\times 10^{7}$ time steps for this system. Nevertheless, our result in Fig. 6 clearly shows an entropy advantage.
@@
As presented in Figure 3, the high entropy advantage is more significant when the neural network is narrower, on the spiral regression task. To further validate this observation, we performed same tests on the other three tasks including Kaggle House Price prediction, MNIST digist recognition, and Polymer SMILES Language modeling. We found the similar trend in all three tasks: narrow networks present larger high-entropy advantage. Detailed results are presented in Figure 7 to 9.
@@
CNN is used in our experiment on the MNIST Digit Recognition task. Therefore, to evaluate the effect of network width on the high entropy advantage, we varied the channel number and tested three different CNNs. Results are presented in Figure 8. While CNNs with 3 and 4 channels present similar level of advantage, the max-entropy state of the 2-channel CNN has significantly higher test accuracy when train loss is small.
@@
We focused on the regressor head of the TransPolymer[33] model in Polymer SMILES language modeling task, which is a linear layer. Therefore, we varied its width from 2 to 100 to evaluate the high entropy advantage at different network width . As we mentioned in the main text, this is a well-learned task with test $R^{2}>0.9$ for all four regressor heads we trained using SGD (see Figure 10). However, max-entropy state is still slightly better than the SGD training state when $W=10$ (Figure 9 (b)). For increased width $W=50$ and $W=100$, the max-entropy state performs equally well with SGD training in at all training losses (Figure 9 (c) and (d)). When we use an extremely narrow network where $W=2$, we can see a significant gap between the max-entropy curve and the SGD training curve. (Figure 9 (a)) While the entropy landscape of in Figure 9 looks different from the others, two independent experiments have been ran to make sure that it is not due to randomness. We speculate this distinct shape of entropy landscape is due to the extreme narrow regressor head employed ($W=2$).

### v2

@@ Discussion
In this paper, we investigated the relationship between the entropy and the generalizability of NNs. We preformed experiments in four distinctive machine learning tasks, covering arithmetic questions, real-world tabular data, computer vision, and language modeling. WLMC and WLMD methods are used to sample the NNs’ entropy landscapes for both classification [$S(L_{train},A_{test})$] and regression tasks [$S(L_{train},L_{test})$]. For classification tasks, at high train loss ($L_{train}$), the max-entropy states achieve random-guess accuracies, supporting the validity of the Wand-Landau-based algorithm in sampling the entropy of NNs; while at low train loss, the max-entropy states generally achieve higher accuracies than the SGD-trained states. For regression tasks, the max-entropy states have lower or at least comparable test losses at all levels of train loss, comparing with their SGD training counterparts. Therefore, for both categories of tasks, the max-entropy states exhibit promising generalizability, defining the high-entropy advantage in neural networks.
Considering entropy is the logarithm of the parameter-space volume, our findings demonstrate that the generalizable well-trained states occupy a larger part of the parameter space than its non-generalizable analog, at the same train loss level. This surprising finding explains why training algorithms can find generalizable states, even in the absence of explicit regularization. Furthermore, max-entropy states frequently surpass the generalization performance of SGD-derived states. This suggests that while SGD (sometimes) favors parameters with marginally lower generalizability, the parameter space possesses a robust and extensive region of generalizable solutions, making the SGD-trained states frequently achieve excellent performance.
@@
The contributions of our work are two-fold. First, such a high-entropy advantage may be widely present in different NN architectures and learning tasks. If so, our work shed lights on a new approach to enhancing model performance without changing its architecture: developing optimizers that can effectively find NNs’ equilibrium (max entropy) states. This could be achieved by methods like low-temperature Newtonian molecular dynamics simulations or simulated annealing. While a few studies have explored this direction before,[42, 43, 44] the existence of the high-entropy advantage in NNs underscores the great potential of further research. Second, the high-entropy advantage also hints at a potential theoretical explanation for the generalizability of overparameterized NNs after SGD training. The SGD training algorithm minimizes the loss function using an estimated gradient of the loss function, which presents strong anisotropy as demonstrated in recent work.[25] The inaccuracy involved in such estimation may bias the trained NN toward not only flat but also higher-entropy minima, which possess higher generalizability.
Our results are also consistent with the philosophical principle of Occam’s razor on a high level, which states that when competing theories of the same observation are presented, simpler ones should be preferred. For an NN with a fixed size, if it encodes a simpler theory of the training data, then fewer neurons are necessary for encoding, and more neurons can vary freely, contributing to a larger entropy. Our finding that these states generalize better provides a concrete reason why theoretical simplicity should be preferred; this also agrees with the observations that regularization can help generalization.
@@
While there are many studies focusing on using machine learning models to solve problems in physics,[45, 46, 47, 48, 49] the reverse directions retains much less attention. However, the foundation concepts of machine learning, especially for neural networks, are deeply rooted in physics. This historic connection is exemplified by the pioneering work including the Hopfield networks[50, 51] and the Boltzmann machine,[52, 53] which is drawn directly from principles of statistical mechanics. We believe our work (re)highlights the potential of treating NNs as physical systems, promoting possibilities to apply well-established statistical physics methods to aid machine learning research.
@@ Methods
Wang-Landau Monte Carlo for Neural Networks
@@
Wand-Landau Monte Carlo (WLMC) is employed for the Spiral Classification Dataset and our results are averaged over 6 independently generated training and testing datasets. For each dataset, 10,000 stages simulations has been performed, where each stage contains $3.2\times 10^{6}$ steps. In each step, we randomly change one parameter in the network and then evaluate both train loss and test accuracy. Therefore, WLMC algorithm has a time complexity of $\mathbb{O}(n)$ to update one trainable parameter, where $n$ is the total number of parameters in the network. We also employed an adapted step size during the simulation by calculating the accepting rate for every 1000 steps. The step size is increased by $10\%$ if the rate is higher than 0.7 and is decreased by $10\%$ when it is lower than 0.3. The “modification factor” in original Wang-Landau method[54] is set to be $f=\exp[5/(i+10)]$ for the i-th stage.
@@
Parameters in the neural networks ranges from $\infty$- ∞ to $\infty$, when no constraint is added. Therefore, the volume of a parameter space is infinite, even for a tiny neural network. On the other hand, to make the thermodynamic quantities valid we must prescribe finite ranges for the parameters. According to early-established experience in neural network training, to alleviate the exploding-gradient and diminishing-gradient problems, the weight parameters should be on the order of $1/\sqrt{W}$.[55] Thus, we limit parameters to the range of $[-2/\sqrt{W},2/\sqrt{W}]$ for our WLMC simulation.
Wang-Landau Molecular Dynamics for Neural Networks
@@
When sampling the entropy landscape of practical neural networks with tens of thousands parameters, WLMC becomes too expensive as it needs to perform the entire forward propagation to update one parameter. Therefore, we employed the Wand-Landau Molecular Dynamics (WLMD) method,[56, 57] which has a time complexity of $\mathbb{O}(n)$ to update all parameters, where $n$ is the total number of trainable parameters in the network. In addition, a Langevin thermostat is implemented to achieve the smooth temperature control during the simulation. The parameters are constrained within the range of $[-2/\sqrt{W},2/\sqrt{W}]$ via reflecting boundary condition. We also find that the range of the parameters does not qualitatively change our results within a reasonable range.
@@
Unlike WLMC, the WLMD algorithm can only directly calculate the entropy as a function of differentiable variables. In order to calculate $S$ as a function of the test accuracy $A_{test}$, we designed a “smoothed” approximation of it:
$A^{s}_{test}=\mbox{Sig}(\alpha\times\delta l),$ (2)
@@
where $\mbox{Sig}(x)=1/[1+\exp(-x)]$ is the sigmoid function, $\alpha$ is a hyperparameter, and $\delta l$ is the difference between the logit of the correct answer and the highest logit of incorrect answers. In the $\alpha\to\infty$ limit, we would have $A^{s}_{test}=A_{test}$. When $\alpha$ is finite, $A^{s}_{test}$ becomes a differentiable approximation of $A_{test}$. In practice we choose $\alpha=5$. This leads to a smoothing error, $|A^{s}_{test}-A_{test}|$, that is always smaller than $1.5\%$ except for the initial $10^{7}$ MD time steps. The initial part of the simulation have significantly higher smoothing errors (as much as $30\%$), but that does not affect the correctness of our final results since WLMD has the capability of starting from an incorrect entropy landscape and gradually converging to the correct entropy landscape.
@@
where $c_{i}=3000$ is a constant and $H(x)=\begin{cases}0,&x<0\\ 1,&x\geq 0\end{cases}$ is the Heaviside step function. Parameters $x_{{max}}$ and $y_{{max}}$ are system-dependent and listed in the supplementary material. By starting with such an $S_{{init}}$, we make the algorithm concentrate its exploration effort in the $x<x_{{max}}$ and $y<y_{{max}}$ regime, improving efficiency. However, the entropy outside this regime becomes invalid, and are thus not presented in this paper.
@@
We verified the correctness of our WLMD protocol by performing traditional (Newtonian) molecular dynamics simulations at a low temperature for the MNIST system. We observed that the test accuracy slowly increases and eventually reaches the high-entropy value found in WLMD simulations. This is consistent with the fact that a system must reach thermodynamic equilibrium in the infinite-time limit. More details are presented in the Supplementary Materials.
@@ Acknowledgements
To verify the correctness of the entropy diagram calculated from Wang-Landau MD simulations, we performed regular (Newtonian) MD of the MNIST system using the train loss as the potential energy. We observed that as the system evolves and approaches equilibrium, the test accuracy indeed increases and approaches the high-entropy value.
The MD simulation was performed at temperature $k_{B}T=0.005$ with Langevin thermostat, and the system eventually equilibrates at potential energy $E_{p}=L_{\mbox{train}}\approx 0.04$. From the entropy diagram (Fig. 2b), at this $L_{\mbox{train}}$ the equilibrium test accuracy is around 0.76, while SGD training achieved a test accuracy of 0.68. Our MD results are presented in Fig. 4. The test accuracy is indeed slowly increasing and approaching 0.76, but the train loss is also slowly decreasing. Since the test accuracy can increase due to either decreasing train loss or increasing entropy (as time passes), we further performed a separation-of-variables study by plotting the test accuracy for only the time steps with the train loss within a small window of $0.0410.04<L_{\mbox{train}}<$ 0.041. For this case, the test accuracy still increases over time, which is an evidence that the test accuracy is indeed positively correlated with entropy.
Figure 4: (a) The train loss (energy) and test accuracy versus time steps (epoches) during Newtonian molecular dynamics simulations of the MNIST system. We averaged over 4 independent instances of simulations to reduce noise. (b) The test accuracy at time steps for which the train loss is in a small range of $0.0410.04<L_{\mbox{train}}<$ 0.041.
@@
Spiral Classification Dataset
We generated the Spiral Classification dataset using the following equations:
@@
where $x_{1i}$ and $x_{2i}$ are the horizontal and vertical coordinates of a point, $r_{i}$ is uniformly distributed between 1 and 5, and $y_{i}$ is the label (i.e., color) of a point. To make it a binary classification task, the label is either 1 (brown) or 0 (green). Both $N_{1i}$ and $N_{2i}$ are Gaussian random noises with mean 0 and standard deviation 0.1.
Figure 5: (a) The Spiral Classification dataset. An independent test set was then generated following exactly the same procedure but with a different random seed. (b) The Spiral Regression dataset. Same as the Classification dataset, an independent test set with 250 points was generated.
@@
We generated Spiral Regression datasets by first generating random coordinates, $-5<x_{1i}<5$ and $-5<x_{2i}<5$, for $N=50$, 100, or 250 points. We then calculate the label (color) of each point using the formula
@@
where $r_{i}$ and $\theta_{i}$ are the polar coordinates of the $i$th point. The training datasets of both classification and regression tasks are presented in Figure 5.
@@
The MNIST handwritten digits dataset is downloaded using the torchvision Python package. We randomly selected 250 images for the training dataset and 250 images for the test dataset. We greatly reduced the size of the dataset in order to increase the difficulty of machine learning and decrease the simulation time.
@@
The Polymer SMILES dataset is collected from the TransPolymer work,[35] which is the foundation chemical language model we used in this study. The dataset is originally built by Ramprasad and coauthors, which consists of $561$ polymer SMILES strings and their bulk electric band gap energy (eV).[58] TransPolymer takes SMILES strings directly into the encoder, which is pretrained over $5$ millions polymer SMILES string via mask language modeling.[35] This pretrained encoder can generate an embedding vector with a dimension of $768$ for each polymer SMILES, which is then fed into the regressor head for downstream task learning. For sampling efficiency, we fixed the encoder of TransPolymer and only finetuning the regressor head, which is also a commonly employed strategy for language model fine-tuning.[39, 40, 35] We follow the same protocol reported in the original TransPolymer work and randomly selected $80\%$ data for train set and the remaining $20\%$ data is used as test set.
@@
To benchmark the max-entropy state, we trained the corresponding neural networks for each task via the SGD optimizer, except for the Sprial Regression task where SGD performs much worse than Adam. For a given network architecture, two hyperparameters have been tuned including learning rate and batch size. Best hyperparamters for different neural networks can be found in Table 1 to 5. Note for Kaggle House Price Prediction task, three different batch sizes have been employed for all neural networks to collect enough data for the full range of train loss. To make sure we are comparing max-entropy state with the best-available SGD training state, we always select the lowest test loss value for SGD training if there are more than one network falls into the bin of a certain train loss level. After determining the best training hyperparameters, we trained $40\sim 200$ neural networks and collected the train loss-test loss (regression task) or the train loss-test accuracy (classification task) trajectories during the entire training process of each task. These trajectories are then used to calculate the SGD (Adam for Sprial Regression) training curves presented in Figure 1 to 3.
@@
To confirm the existence of entropy advantage in deeper neural networks with more complex architectures, we also constructed a 10-layer ResNet.[11] The neural network consists of an initial convolutional layer with 6 channels, and then 4 residual blocks with 6, 12, 12, and 12 channels each, and a final fully-connected layer. The original ResNet architecture contains batch-normalization operations, which breaks the rigorous analogy between training losses and physical energy landscapes. Therefore, we replaced all of them with layer normalization operations. Our training and test datasets each contain 2500 randomly chosen images of either a deer or a frog from the CIFAR-10 dataset. Due to the significantly higher computational cost, we only ran WLMD for $3\times 10^{7}$ time steps for this system. Nevertheless, our result in Fig. 6 clearly shows an entropy advantage.
@@
As presented in Figure 3, the high entropy advantage is more significant when the neural network is narrower, on the spiral regression task. To further validate this observation, we performed same tests on the other three tasks including Kaggle House Price prediction, MNIST digit recognition, and Polymer SMILES Language modeling. We found the similar trend in all three tasks: narrow networks present larger high-entropy advantage. Detailed results are presented in Figure 7 to 9.
@@
CNN is used in our experiment on the MNIST Digit Recognition task. Therefore, to evaluate the effect of network width on the high entropy advantage, we varied the channel number and tested three different CNNs. Results are presented in Figure 8. While CNNs with 3 and 4 channels present similar level of advantage, the max-entropy state of the 2-channel CNN has clearly higher test accuracy when train loss is small.
@@
We focused on the regressor head of the TransPolymer[35] model in Polymer SMILES language modeling task, which is a linear layer. Therefore, we varied its width from 2 to 100 to evaluate the high entropy advantage at different network width . As we mentioned in the main text, this task is reported to be well-learned. [35] Our results further verify this as the test $R^{2}$ exceeded 0.9 for all four regressor heads we trained using SGD (see Figure 10). However, max-entropy state is still slightly better than the SGD training state when $W=10$ (Figure 9 (b)). For increased width $W=50$ and $W=100$, the max-entropy state performs equally well with SGD training in at all training losses (Figure 9 (c) and (d)). When we use an extremely narrow network where $W=2$, we can see a larger gap between the max-entropy curve and the SGD training curve. (Figure 9 (a)) While the entropy landscape of in Figure 9 (a) looks different from the others, two independent experiments have been ran to make sure that it is not due to randomness. We speculate this distinct shape of entropy landscape is due to the extreme narrow regressor head employed ($W=2$).